OpenStudy (anonymous):

um x+2/5(x-3) divide x-2/5(x-3)

OpenStudy (anonymous):

Flip n multiply the second fraction. The 5(x-3) cancel. x+2/(x-2)

OpenStudy (anonymous):

can you cancel when there is no equal sign?

jimthompson5910 (jim_thompson5910):

Yes, x/x = 1 (ie x/x simplifies to 1 even though it's not explicitly equal to anything

jimthompson5910 (jim_thompson5910):

So what I'm saying is that you can simplify even when there is no equal sign.

OpenStudy (anonymous):

Well you can leave it there if you want, it's the same equation, just not simplified. The top and bottom are being multiplied by the same number so it's unnecessary.

OpenStudy (anonymous):

so what your saying is that ^ is = x+2/5(x-3) * 5(x-3)/x-2?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

If you're dividing fractions you can flip the second one and multiply for ease of calculating :)

OpenStudy (anonymous):

(x+2)/(x-2)

OpenStudy (anonymous):

umm restriction is x cannot be 2 u have to state that else its wrong

OpenStudy (anonymous):

er, multiply top by bottom?

OpenStudy (anonymous):

wtf so confused lol

OpenStudy (anonymous):

umm give me a sec

OpenStudy (anonymous):

ah crap i dont know hw to use eqn thing but u flip when dividing x+2 5(x-3) ------- x -------- 5(x-3) x-2

OpenStudy (anonymous):

do u see it

OpenStudy (anonymous):

mhmm but wouldn't that be (x+2)(x-2) instead of (x+2)/(x-2)?

OpenStudy (anonymous):

how would u conclude that?

OpenStudy (anonymous):

er... I don't know lol.

OpenStudy (anonymous):

25(x-3)²/(x-2)(x+2)

OpenStudy (anonymous):

lol no i already flipped it, just cancel the 5(x-3) out and restriction is x cannot be 3 or 2

OpenStudy (anonymous):

x+2/5(x-3) * 5(x-3)/(x-2) gives (x+2)*5(x-3) / (5(x-3)*(x-2)) The 5(x-3)'s cancel, you're left with... (x+2)/(x-2)

OpenStudy (anonymous):

facepalm. what's this buisness with the restrictions?

OpenStudy (anonymous):

It won't exist when the denominator = 0

jimthompson5910 (jim_thompson5910):

that's to avoid division by zero

OpenStudy (anonymous):

well u cannot divide by 0

OpenStudy (anonymous):

so x can't = 2.. only?

jimthompson5910 (jim_thompson5910):

and can't equal 3 as well

OpenStudy (anonymous):

what? why? we're left with just (x+2)/(x-2)

OpenStudy (anonymous):

Why 3?

jimthompson5910 (jim_thompson5910):

because we need to ensure that the final expression is equal to the original expression

jimthompson5910 (jim_thompson5910):

it's like saying f(x) = g(x), which would imply that f and g have the same domain

OpenStudy (anonymous):

i sort of get it. (x-3) and all

OpenStudy (anonymous):

mybrain won this round though.

OpenStudy (anonymous):

mybrain 1 me 0

jimthompson5910 (jim_thompson5910):

basically, to find the restrictions, look at the original expression, not the final one

jimthompson5910 (jim_thompson5910):

Back to the basic example of x/x. Clearly x/x=1, but when x=0, x/x is undefined (it's actually an indeterminate expression, you'll learn about this later), but 1 is clearly defined. So something went terribly wrong here. So to avoid this, we just say that x can't equal zero.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

thanks guys, I'll have my girlfriend's dad explain this to me in real life.

jimthompson5910 (jim_thompson5910):

just remember that something like x=x is only true if EVERY value of x works (ie makes that equation true). If a very select few don't work, then you have to exclude them

OpenStudy (anonymous):

But haven't we simplified so we can solve the equation when x = 3. I can tell you it will equal 5 but you're saying that's incorrect, because if we have the complete equation it would be indeterminable. Also there is nothing that says i can't multiply (x-1)/(x-1) by the equation. It will be the same equation but now you also can't calculated it when x = 1.

OpenStudy (anonymous):

So by that same logic, every equation is indeterminable at all values of x.

jimthompson5910 (jim_thompson5910):

All I'm doing is looking at the original expression. Since there is an x-3 in the denominator, plugging in x=3 will result in a division by zero error. So this is why x cannot equal 3.

OpenStudy (anonymous):

Or the orginal equation is absolute and simplifying it will result in a loss of information?

jimthompson5910 (jim_thompson5910):

Yes that's one way of putting it. You "lose" the info that x=3 is not in the domain. So if that's a crucial or critical value, you've essentially lost it in translation.

OpenStudy (anonymous):

But i could represent that equation as (x+n)/(x+n)*(x+2)/(x-2) with n = -infinity to +infinity and then it can not be calculated for any value of x. Which i know is not the case, right?

jimthompson5910 (jim_thompson5910):

That may be true, but because you're introducing these discontinuities, I don't think they affect the original expression's domain. You could take your argument even further and take something really simple like x and blow it up to get x*(x+m)/(x+m)*(x-n)/(x-n)*... and you could go on forever introducing discontinuities, but clearly x can be any number.